OpenStudy (anonymous):

How to I find the two values of theta in degrees (0 deg <_ theta <_ 360 deg) for tan theta = sqrt 3 using the unit circle?

OpenStudy (anonymous):

p.s. i know how to find sine and cosine, just need help finding tangent! :)

OpenStudy (fibonaccichick666):

What can you tell me about special angles?, can you draw the unit circle?

OpenStudy (fibonaccichick666):

Ah, ok so how do you express tangent in terms of sine and cosine?

OpenStudy (anonymous):

i have a picture of it right in front of me, i just don't know what i should be looking for

OpenStudy (anonymous):

would tangent represent x?

OpenStudy (anonymous):

\[\tan(\theta)=\sqrt3\] means \[\frac{\sin(\theta)}{\cos(\theta)}=\sqrt3 \]which will occur if \[\sin(\theta)=\frac{\sqrt3}{2}, \cos(\theta)=\frac{1}{2}\] because \[\frac{\frac{\sqrt3}{2}}{\frac{1}{2}}=\sqrt3\]

OpenStudy (fibonaccichick666):

no, so here is an important identity. Tan x= (sin x) / (cos x)

OpenStudy (anonymous):

would that be 5pi/6 pi/6 and pi/3?

OpenStudy (fibonaccichick666):

think about the value of sine cosine and tangent, in which quadrant are they all positive?

OpenStudy (anonymous):

in quadrant 1

OpenStudy (anonymous):

so the two values would just be pi/3 and pi/6 @FibonacciChick666 ??

OpenStudy (fibonaccichick666):

You should only have one answer, check their sines and cosines only one yields the correct answer

OpenStudy (anonymous):

but the problem says to find two values

OpenStudy (fibonaccichick666):

well, one of those

OpenStudy (fibonaccichick666):

there is another quadrant where the tangent would be positive

OpenStudy (fibonaccichick666):

hence your other answer